solve the quadratic equations for the value of x.
X^2+x=12

Answer by ewatrrr(24785)

 
Hi,
Solving for x
x^2+x = 12
x^2 + x - 12 = 0
factoring
(x+4)(x-3)= 0
(x+4)= 0 x = -4
(x-3)= 0 x = 3